7).

a.Let the sample space be S = {1, 2, 3, 4, 5, 6, 7, 8, 9, 10}. Suppose the outcomes are equally likely. Compute the probabilityb.The data in the following table show the association between cigar smoking and death from cancer for 139,625 men. Note: curre
Answer #1

a)

Given

Sample space , S = {1,2,3,4,5,6,7,8,9,10}

Total number of elements in the sample space = 10

Define E = " an odd number less than 7"

Odd number less than 7 = {1,3,5}

Number of odd number less than 7 = 3

Now

P(E) = 3 / 10 = 0.3

b)

Died from Cancer Did not die from Cancer Total
Never smoked cigars 676 121252 121928
Former cigar smoker 65 8726 8791
Current cigar smoker 131 8775 8906
Total 872 138753 139625

a.

Total number of men = 139625

Number of individuals Died from Cancer = 872

P( died from cancer) = 872 / 139625 = 0.006245 \approx 0.006

The probability that he died from cancer = 0.006

b.

Number of individuals current cigar smoker = 131 + 8775 = 8906

P(current cigar smoker ) = 8906 / 139625 = 0.063785  \approx 0.064

The probability that he was a current cigar smoker = 0.064

c.

Number of individuals that died from cancer and was a current cigar smoker = 131

P( died from cancer and current cigar smoker ) = 131 / 139625 = 0.000938 \approx 0.001

The probability that he died from cancer and was a current cigar smoker = 0.001

d.

From above calculation;

P(died from cancer) = 0.006

P(current cigar smoker ) = 0.064

P(died from cancer and current cigar smoker )= 0.001

Now,

P( died from cancer or current cigar smoker )

= P(died from cancer) + P(current cigar smoker ) - P(died from cancer and current cigar smoker )

= 0.006 + 0.064 - 0.001

= 0.069

The probability that he died from cancer or was a current cigar smoker = 0.069